### v1

# Designing AI-Resilient Assessments Using Interconnected Problems: A Theoretically Grounded and Empirically Validated Framework

Abstract
The proliferation of generative AI has rendered traditional, modular assessments in computing education obsolete, creating a critical disconnect between academic training and industry practice. This paper presents a theoretically grounded framework for designing AI-resilient assessments, supported by formal proofs and empirical validation. We make three primary contributions. First, we establish two formal theorems: (1) assessments composed of interconnected problems, where outputs serve as inputs to subsequent stages, are inherently more AI-resilient than modular assessments due to multi-step reasoning and context limitations of large language models; and (2) semi-structured problems with deterministic success criteria provide more reliable measures of student competency than fully open-ended projects, which allow AI systems to default to familiar solution patterns. These findings challenge widespread recommendations from international policy organizations [unesco2025assessment] and institutional guidelines [cornell2024ai, wustl2024ai] that open-ended, authentic assessments are the primary solution to maintaining academic integrity in the AI era. Second, we validate these theorems through empirical analysis of four university data science courses ($N=138$), demonstrating a substantial AI inflation effect: students achieve near-perfect scores (9090100100%) on AI-assisted modular homework but drop approximately $30$ percentage points on proctored exams (Cohen’s $d=1.51$). Critically, interconnected projects maintain strong alignment with modular assessments ($r=0.954$, $p<.001$) while remaining AI-resilient, whereas proctored exams show only moderate correlation ($r=0.726$, $p<.001$), suggesting they assess different competencies. Third, we operationalize these findings into a practical design procedure that translates theoretical principles into implementable strategies, providing computing educators with a systematic, evidence-based approach for designing assessments that encourage deeper engagement, authentically reflect industry practice, and naturally resist trivial AI delegation.
I Introduction
Since the release of generative AI tools like ChatGPT in late 2022, the rapid adoption of this technology has reshaped professional practice at a fundamental level [mckinsey2024state, ey2024work]. Tasks that once defined early-career technical work—drafting boilerplate code, writing simple SQL queries, or executing small-scale data analysis—are now completed instantly by AI tools, a shift confirmed by large-scale industry experiments [dellacqua2023navigating]. Yet, these same modular tasks continue to dominate academic assessments, creating a widening and unsustainable gap between university curricula and actual industry requirements [jin2025generative]. This misalignment raises a critical question: What is the educational value of practicing skills that industry has already automated? Students quickly recognize this disconnect, undermining their motivation and learning when assignments feel obsolete [dempere2023impact].
While educators widely acknowledge this gap, proposed solutions often focus on how instructors can use AI as a teaching assistant—for intelligent tutoring, adaptive feedback, or automated grading—rather than on how assessment design itself must fundamentally change [ZawackiRichter2019, BondEtAl2024]. Such an approach overlooks the core challenge: the nature of meaningful technical work has shifted. Professionals now focus less on executing discrete tasks and more on orchestrating complex, multi-step workflows that integrate AI as a component, a form of higher-order work that current AI systems cannot replicate [brynjolfsson2023macroeconomics]. This shift demands a new pedagogical focus on AI-resilient assessments: assignments that cultivate genuine learning precisely because they cannot be fully delegated to an AI.
Our approach is grounded in the recognized limitations of large language models (LLMs). While proficient at single-step, modular problems, even state-of-the-art models like GPT-$5$ exhibit brittleness in tasks requiring sustained, multi-step reasoning [phan2025humanity, shojaee2025illusion]. Empirical benchmarks consistently show that LLMs struggle with long inference chains, logical negation, and discovering abstract structural patterns, often exploiting superficial token biases rather than engaging in genuine inference [parmar2024logicbench, gendron2024abstractreasoners, jiang2024tokenbias]. This limitation is not merely theoretical; domain-specific evaluations in fields like biostatistics confirm that LLMs fail on multi-stage procedures unless guided by iterative human feedback [ignjatovic2023biostat]. The boundary of AI capability, therefore, defines the frontier of essential human skill.
This paper argues that the key to AI-resilient assessment lies in designing projects with a high degree of interconnectedness. Based on empirical observations from teaching four sessions of upper-level courses in data science and statistical learning, we make two central claims. First, we contend that complexity, defined by the interdependence among a project’s components, is the primary mechanism for fostering the integrative problem-solving skills that current AI tools lack and that industry now demands. Second, we challenge the widespread recommendation—articulated in international policy documents [unesco2025assessment] and institutional guidelines [cornell2024ai, wustl2024ai]—that open-ended, authentic assessments are the primary solution to maintaining academic integrity in the age of generative AI. While these recommendations rest on the assumption that AI ‘struggles with original thought and complex synthesis’ [unesco2025assessment], recent evidence demonstrates that LLMs’ limitations stem not from an inability to generate creative responses, but from fundamental architectural constraints in multi-step reasoning and context management [phan2025humanity, shojaee2025illusion].
II Related Work
Our work is situated at the intersection of three distinct but related research areas: ($1$) the impact of generative AI on academic integrity and assessment, including recent evidence that generative AI can harm learning when used without appropriate guardrails [bastani2025generative, Cotton2023, Sullivan2023, Yan2024]; ($2$) the principles of authentic assessment and complex problem-solving in STEM education [Ullah2020, Csapo2017, Herde2016]; and ($3$) the cognitive science of task complexity [Chen2023]. While no prior work directly proposes a framework for AI-resilience based on interconnectedness, these adjacent fields establish the foundation and necessity for our contribution.
II-A Generative AI, Assessment, and Academic Integrity
The release of ChatGPT precipitated a surge of academic literature focused on the challenges generative AI poses to traditional assessment and academic integrity. A significant body of this work documents the scope of the problem, highlighting the ease with which students can use AI to complete assignments, thereby undermining learning and creating an integrity crisis [Cotton2023, Sullivan2023, BinNashwan2023]. These studies effectively diagnose the issue, arguing that the availability of powerful AI necessitates a fundamental rethinking of assessment practices. However, the proposed solutions often remain high-level, advocating for a shift towards in-class assessments, oral examinations, or a greater emphasis on process over product, without offering a systematic, scalable framework for redesigning coursework itself [Yan2024].
Moreover, overly focusing on in-class assessments may test a completely different skill—test-taking ability rather than the intended learning outcomes [bastani2025generative]. Forgoing project-based assessment deprives students of the opportunity to practice skills that closely mimic real-world project work by industry professionals [Ullah2020, argus2022embedding]. We argue that adding guardrails through designing AI-resilient assessments is more beneficial, as it preserves authentic learning experiences while addressing integrity concerns.
II-B Authentic Assessment and Complex Problem-Solving
Our approach builds on a long tradition of research into authentic assessment and complex problem-solving. Authentic assessment emphasizes mirroring the challenges and standards of professional practice, a principle well-established in engineering and computing education [Ullah2020]. Similarly, the field of complex problem-solving (CPS) focuses on assessing cognitive processes in dynamic, non-routine situations, which aligns with our goal of moving beyond static, modular tasks [Csapo2017, Herde2016]. While philosophically aligned, this body of work predates the generative AI era and thus does not explicitly address how to design assessments that are resilient to AI delegation. It provides the pedagogical why—the need for real-world problem-solving—but not the operational how in a world where AI can solve many seemingly complex but isolated tasks.
II-C Task Complexity and Cognitive Load Theory
To ground our concept of interconnectedness, we draw on Cognitive Load Theory (CLT), particularly the concept of element interactivity. CLT posits that the intrinsic difficulty of a task is determined by the number of interacting elements that must be simultaneously processed in working memory [Chen2023]. A task with low element interactivity (e.g., solving a series of independent coding problems) can be broken down and solved piece-by-piece. In contrast, a task with high element interactivity (e.g., debugging a system where a change in one module has cascading effects on others) requires a holistic understanding of the entire system. We argue that current generative AI excels at tasks with low element interactivity but struggles with high-interactivity problems. Therefore, designing assessments with high element interactivity—what we term interconnectedness—is a theoretically grounded approach to creating AI-resilient learning experiences. Building on this foundation, we now formalize these principles through two theorems that establish the theoretical basis for AI-resilient assessment design.
III Methodology: A Framework for AI-Resilient Assessment
Our methodology is grounded in a theoretical framework that we first establish through two theorems. We then proceed to validate these theorems through a multi-year empirical study.
III-A Theoretical Framework
We begin by formally stating and proving two theorems that underpin our assessment design principles.
Theorem 1.
An assessment composed of interconnected problems, where the output of one problem serves as the input for the next, is more AI-resilient than an assessment composed of equivalent but modular (independent) problems.
That is, chaining problems together increases the difficulty for a generative AI model in a way that solving them independently does not, because it forces multi-step reasoning and strains the model’s limited memory (context window).
Proof of Theorem 1.
Let an assessment be a sequence of problems $P=(p_{1},p_{2},...,p_{n})$. In a modular design, each $p_{i}$ is independent. In an interconnected design, the solution to $p_{i+1}$ is contingent upon the solution of $p_{i}$. A generative AI model must therefore solve the entire sequence $(p_{1},...,p_{i})$ to correctly solve $p_{i}$. This requirement increases the computational and reasoning load on the model in two primary ways. First, it necessitates multi-step, sequential reasoning, a known weakness of current large language models (LLMs) where performance degrades as the number of steps increases [Wei2022, Kojima2022]. Second, the cumulative context required to solve later problems in the sequence can exceed the model’s context window, leading to information loss and failure [Narayanan2023]. Because a modular design allows each problem to be solved in a separate, stateless context, it avoids these vulnerabilities. Thus, an interconnected design is inherently more resilient to trivial AI delegation and better reflects the complex, stateful workflows common in professional practice [Brynjolfsson2023]. ∎
Theorem 2.
A semi-open-ended project, which provides a clear structure and deterministic success criteria, is a more reliable measure of student competency than a fully open-ended project when generative AI tools are ubiquitous.
In other words, by giving a specific problem with a clear goal, we prevent the AI from choosing an easier, more familiar problem for which it has been heavily trained, which gives us a truer measure of the student’s ability to solve the actual task.
Proof of Theorem 2.
Large language models (LLMs) are fundamentally trained to minimize perplexity on their training distribution, which biases them toward generating sequences that closely match patterns encountered during pre-training [Huang2025Verbatim, Carlini2022Quantifying]. When presented with a fully open-ended problem (e.g., “analyze a dataset of your choice”), the unconstrained (open-ended problem) solution space allows the model to select a problem instance that lies within the model’s training distribution, where memorized patterns and dataset artifacts can be exploited as shortcuts [Du2022Shortcut]. This phenomenon, known as shortcut learning, occurs when models rely on spurious correlations rather than robust reasoning, significantly degrading generalizability to out-of-distribution tasks [Du2022Shortcut, Tang2023Lazy].
Recent work has demonstrated that LLMs exhibit lazy learning behavior, preferentially following the path of least resistance when afforded flexibility in problem selection [Tang2023Lazy, Sanyal2025Path]. Specifically, Tang et al. [Tang2023Lazy] show that larger models are more likely to exploit shortcuts during inference, while Sanyal et al. [Sanyal2025Path] demonstrate that policy optimization in LLMs “consistently follows the path of least resistance,” avoiding complex reasoning when simpler alternatives are available. In the context of open-ended assessments, this translates to students (with AI assistance) selecting problem variants with low perplexity—those that align closely with the model’s training data—thereby producing high-quality outputs that do not reflect genuine problem-solving ability on novel or challenging tasks.
Furthermore, research on verbatim memorization reveals that sequences with lower perplexity are significantly more likely to be memorized and reproduced by LLMs [Huang2025Verbatim]. Huang et al. demonstrate that “sequences that are better recognized by the language model, i.e., sequences with lower perplexity, are more likely to be verbatim memorized,” and that this memorization is “intertwined with the LM’s general capabilities” [Huang2025Verbatim]. Consequently, when students are permitted to choose their own datasets or problem instances, they (or the AI tools they employ) will gravitate toward familiar, well-represented problem types in the training corpus, where the model can leverage memorized solution templates rather than engaging in novel reasoning.
A semi-open-ended project addresses this limitation by constraining the solution space through two mechanisms: (1) specifying a particular dataset or problem instance, and (2) defining deterministic success criteria (e.g., “achieve $90\%$ F1-score on this held-out test set”). This design forces engagement with a potentially out-of-distribution task, where the model cannot simply retrieve memorized solutions or exploit dataset artifacts. The deterministic evaluation metric further prevents the model from optimizing for plausibility (a subjective, perplexity-driven objective) rather than correctness (an objective, task-specific objective). By narrowing the optimization landscape to a specific, well-defined problem, semi-open-ended assessments compel students to demonstrate competency on the assigned task, yielding a more reliable and robust measure of their problem-solving ability [Ullah2020].
In summary, the adversarial robustness of semi-open-ended assessments stems from their resistance to the fundamental optimization behavior of LLMs: the tendency to exploit shortcuts, memorize training data patterns, and follow paths of least resistance. By constraining the problem space and imposing deterministic evaluation, we force both students and AI tools to engage with the specific complexities of the task, providing a more authentic measure of competency in an AI-saturated educational environment. ∎
III-B Empirical Validation
To validate these theorems, we analyzed assessment data from four offerings of upper-level undergraduate and graduate courses in data science spanning multiple academic years. This section details the course context, the assessment designs, and the data analysis procedures.
III-B1 Study Context and Population
The study was conducted across four course offerings at a large research university between Fall 2024 and Spring 2025:
  • Predictive Analytics (Fall 2024): An undergraduate course (N=34) with a mix of modular assignments and an open-ended final project.
  • Data Mining (Fall 2024): An undergraduate course (N=41) also with a mix of modular assignments and an open-ended final project.
  • Python for Data Science and AI (Spring 2025): A graduate-level course (N=41) featuring a mix of modular assignments, interconnected projects, and proctored exams.
  • Predictive Analytics (Fall 2025): The same undergraduate course (N=22) but with a mix of modular assignments and a redesigned, interconnected final project.
III-B2 Data Collection and Analysis
For each course, we collected anonymized student grade data for all assessments. Our analysis proceeds in three stages:
  1. 1.
    Descriptive Analysis: We calculate descriptive statistics for each assessment type.
  2. 2.
    Comparative Analysis: We use paired-samples t-tests to compare student performance on modular vs. project-based assessments and independent-samples t-tests to compare open-ended vs. interconnected projects. Effect sizes (Cohen’s d) are calculated.
  3. 3.
    Correlation Analysis: We use Pearson correlations to examine the relationship between performance on different assessment types.
All analyses were conducted using Python with the SciPy and statsmodels libraries. Visualizations were generated using Matplotlib and Seaborn.
IV Results
IV-A The AI Inflation Effect: Modular vs. Proctored Assessments
In the Spring 2025 Python for Data Science and AI course (N=41), we observed a stark contrast in student performance between modular assessments (knowledge checks and homework, where AI use was permitted) and proctored exams (where AI use was prohibited). As shown in Figure 1, students achieved near-perfect scores on knowledge checks (M=92.55, SD=20.26), but their performance dropped significantly on the proctored exams (M=62.82, SD=18.49). This represents a performance gap of nearly 30 percentage points, with a large effect size (Cohen’s d = 1.52).
Figure 1: Performance on Different Assessment Types in Python for Data Science and AI (Spring 2025).
Table I provides a detailed breakdown of this AI inflation effect, including statistical significance and effect sizes for each assessment type compared to the proctored exam baseline. More tellingly, while 90.2% of students scored 90% or higher on knowledge checks, only 4.9% achieved this on the exams. This dramatic shift in the distribution of high performers suggests that the high scores on modular assignments may be inflated by AI assistance and do not reflect true mastery of the material. The consistency of this effect across both knowledge checks and homework (Cohen’s d = 1.54 and 1.48, respectively) indicates that the AI inflation phenomenon is robust across different types of modular assessments.
TABLE I: AI Inflation Effect: Performance Gap Between AI-Permissive and Proctored Assessments in Python for Data Science (Spring 2025, N=41)
Assessment Type Mean (SD) % $\geq$ 90 $\Delta$ from Exam Cohen’s d
Knowledge Checks 92.55 (20.26) 90.2% +29.73*** 1.54
Homework 88.73 (15.84) 70.7% +25.91*** 1.48
Modular Average 90.64 (16.89) 80.5% +27.82*** 1.52
Proctored Exams 62.82 (18.49) 4.9%
*** $p<0.001$ (paired t-test)
IV-B Interconnectedness as a Differentiator: Open-Ended vs. Interconnected Projects
Our comparison of the Fall 2024 Predictive Analytics course (open-ended project) and the Fall 2025 offering (interconnected semi-open-ended project) reveals a more nuanced picture. While the average project scores were not dramatically different (91.47 vs. 89.66), the interconnected design led to a greater performance drop from the modular assignments and, critically, increased score variability. Figure 2 illustrates this comparison through box plots showing the distribution of scores.
Figure 2: Comparison of Performance Drop from Modular to Project Assessments. The interconnected project design shows a greater (though not statistically significant) drop and higher variability.
The performance drop from the modular average to the final project was larger in the interconnected course (3.07 points) than in the open-ended course (1.82 points). More importantly, the standard deviation of project scores was higher in the interconnected course (SD=21.93) compared to the open-ended course (SD=16.83), representing a 30% increase in absolute variability. This increased variance is not a flaw; it is evidence that the interconnected design created a more challenging and authentic assessment that could not be trivially solved by invoking an AI for each sub-problem. Table II provides a comprehensive overview of assessment performance across all four course offerings, highlighting the consistent patterns of performance gaps and correlations.
TABLE II: Summary of Assessment Performance Across Four Course Offerings
Course N Modular Project/Exam Performance Correlation Effect
Mean (SD) Mean (SD) Gap (r) Size (d)
Courses with Open-Ended Projects (Fall 2024)
     Predictive Analytics 34 93.29 (12.45) 91.47 (16.83) 1.82-1.82 0.89*** 0.12
     Data Mining 41 94.15 (8.92) 92.68 (14.25) 1.47-1.47 0.76*** 0.13
Course with Interconnected Project and Proctored Exam (Spring 2025)
     Python for Data Science and AI 41 90.64 (16.89) 62.82 (18.49) 27.82-27.82*** 0.73*** 1.52
Course with Interconnected Project (Fall 2025)
     Predictive Analytics 22 92.73 (15.67) 89.66 (21.93) 3.07-3.07 0.91*** 0.16
*** $p<0.001$; Performance Gap = Project/Exam Mean - Modular Mean
Modular = average of homework and knowledge checks; Effect Size = Cohen’s $d$
IV-C Predictive Validity of Assessments
We examined how well performance on modular assignments predicted performance on other assessment types. Table III presents a comprehensive correlation matrix for the Python course, revealing distinct patterns in how different assessment types relate to one another.
TABLE III: Correlation Matrix: Relationships Between Assessment Types in Python for Data Science Course (N=41)
Assessment Type Knowledge Homework Project Exam
Checks
Knowledge Checks 1.000
Homework 0.861*** 1.000
Project (Interconnected) 0.936*** 0.882*** 1.000
Proctored Exam 0.725*** 0.603*** 0.752*** 1.000
Modular Average 0.964*** 0.958*** 0.954*** 0.726***
*** $p<0.001$
Modular Average = (Knowledge Checks + Homework) / 2
The correlation between the modular average and exam scores was moderate (r = 0.726, p < .001), while the correlation between the modular average and the interconnected project scores was much stronger (r = 0.954, p < .001). This pattern is visualized in Figure 3, which presents scatter plots comparing these relationships. Panel (a) shows the tight clustering of points around the regression line for modular-project correlation, while panel (b) shows greater dispersion for modular-exam correlation.
Figure 3: Comparison of Predictive Validity: (a) Modular assessments strongly predict interconnected project performance (r=0.954), indicating both measure similar skills under different conditions. (b) Modular assessments moderately predict exam performance (r=0.726), suggesting exams may measure different skills or are less susceptible to AI inflation.
This finding suggests that interconnected projects, while more challenging, are a better measure of the same underlying skills being taught in the modular assignments. The strong correlation (r=0.954) indicates construct validity: both assessment types are measuring the same competencies, but the interconnected design does so under conditions that are less susceptible to AI inflation. In contrast, the moderate correlation with exams (r=0.726) may indicate that proctored exams are either testing a different set of skills (e.g., memorization, speed under pressure) or are simply less susceptible to AI influence due to their controlled environment. The visual contrast in Figure 3 makes this distinction immediately apparent, showing that interconnected projects preserve the pedagogical alignment with course content while adding AI-resilience.
V Discussion
Our findings provide empirical support for the two theorems established in our theoretical framework. The 30-point performance gap between AI-permissive modular assignments and proctored exams (Table I) demonstrates what we term an AI inflation effect, where assessment scores may not accurately reflect student mastery when AI tools are available. While proctored exams are less susceptible to AI inflation, they may assess different competencies (e.g., time pressure, memorization) rather than the intended learning outcomes. Our data indicate that semi-structured, interconnected projects offer a third path: they preserve the authentic, workflow-based learning experiences that mirror industry practice while providing better differentiation of student ability than fully open-ended designs when AI tools are available.
More importantly, our results demonstrate that interconnectedness is a useful mechanism for creating AI-resilient assessments. Comparing two offerings of the same Predictive Analytics course—one with an open-ended project (Fall 2024, N=34) and one with an interconnected project (Fall 2025, N=22)—the interconnected design exhibited higher score variability (SD=21.93 vs. 16.83) despite similar mean scores (89.66 vs. 91.47, Table II). This increased variance is not a flaw; it is a feature. It indicates that the interconnected project was a better differentiator of student ability, creating a more challenging and authentic assessment that could not be trivially solved by invoking an AI for each sub-problem. This supports Theorem 1, which posits that forcing multi-step, stateful reasoning increases resilience to AI delegation.
Furthermore, the strong correlation between modular assessments and interconnected project scores (r=0.954, p<.001, Table III) suggests that both are measuring similar underlying skills. However, the interconnected design does so under conditions that are less susceptible to AI inflation, making it a more robust measure of those skills. In contrast, the moderate correlation between modular assessments and proctored exam scores (r=0.726, p<.001, Table III) suggests that exams may measure different competencies or are influenced by factors such as time pressure and test-taking ability. This addresses the core challenge of modern assessment: how to evaluate the skills taught in a curriculum in a way that is both authentic and AI-resilient.
Our findings also lend support to Theorem 2. The high scores and low variance on open-ended projects suggest that students (likely with AI assistance) may be choosing a path of least resistance, solving a simpler version of the problem than intended. The semi-open-ended, interconnected design, by contrast, constrains the problem space and forces engagement with the specific, intended challenges, resulting in a more reliable measure of competency.
V-A A Design Procedure for Implementation
To translate these theoretical insights into actionable guidance for educators, we present a systematic design procedure for AI-resilient assessments (Algorithm 1). This procedure operationalizes Theorems 1 and 2 through six concrete steps that can be adapted to various computing education contexts.
Algorithm 1 Design Procedure for AI-Resilient Assessments
0: Learning objectives $L$, course content $C$, target competencies $T$
0: AI-resilient assessment design AA
1:Step 1: Identify Core Workflow: $W\leftarrow$ Map $T$ to multi-stage workflow $\triangleright$ e.g., clean $\rightarrow$ EDA $\rightarrow$ model $\rightarrow$ evaluate
2:Step 2: Create Sequential Dependencies (Theorem 1): $stages\leftarrow$ Decompose $W$ where $input(s_{i})=output(s_{i-1})$
3:Step 3: Impose Semi-Structured Constraints (Theorem 2): Define deterministic success criteria $\triangleright$ e.g., F1 >0.90>0.90 on test set
4:Step 4: Embed Verifiable Checkpoints: For each stage, define intermediate output with validation $\triangleright$ e.g., zero nulls
5:Step 5: Introduce Domain Complexity: Embed challenges requiring contextual reasoning $\triangleright$ e.g., class imbalance
6:Step 6: Validate AI-Resilience: Pilot with cohort; measure $r_{\textrm{project, other}}$, $\sigma_{\textrm{project}}$, $\textrm{time}_{\textrm{median}}$
7:return AA $\triangleright$ AI-resilient assessment design
The algorithm emphasizes several key principles. First, Step 2 directly operationalizes Theorem 1 by creating sequential dependencies that force multi-step reasoning and exceed typical AI context windows. Second, Step 3 implements Theorem 2 by constraining the solution space through semi-structured requirements, preventing AI from defaulting to easier problem variants. Third, Steps 4 and 5 introduce additional layers of complexity—verifiable checkpoints and domain-specific challenges—that further increase AI-resilience while maintaining pedagogical validity.
The validation step (Step 6) is particularly critical. Our empirical results suggest that a well-designed interconnected project should exhibit three characteristics: (1) a strong correlation with other assessments of the same skills, indicating construct validity; (2) higher score variance compared to modular or open-ended alternatives, indicating better differentiation; and (3) substantial student time investment, indicating genuine engagement. These metrics provide instructors with concrete criteria for evaluating whether their redesigned assessments are achieving the intended goals.
VI Conclusion and Future Work
This paper makes three primary contributions to the field of computing education in the age of generative AI. First, we establish a theoretical foundation through two formal theorems that position interconnectedness and semi-structured design as key principles for creating AI-resilient assessments. Second, we validate this framework with empirical evidence from four university data science courses, demonstrating a substantial AI inflation effect in traditional modular assessments while showing that interconnected projects preserve construct validity and provide superior differentiation of student ability. Third, we operationalize these findings into a practical design procedure that translates theoretical principles into implementable assessment strategies.
Our work has immediate practical implications for educators. By shifting focus from the difficulty of individual tasks to the complexity of the connections between them, instructors can design assignments that preserve authentic, workflow-based learning experiences while naturally resisting trivial AI delegation. The framework we propose offers a systematic, theoretically-grounded alternative to ad-hoc solutions like banning AI or reverting to traditional in-person exams, which may assess different competencies than the intended learning outcomes.
Future work will aim to replicate these findings across a broader range of institutions and disciplines to establish the generalizability of our framework. Further research could also explore the optimal level of interconnectedness, as it is likely that there is a point of diminishing returns where cognitive load becomes excessive. Finally, the development of automated tools to help instructors design and evaluate interconnected assessments would be a valuable contribution to the field.